OpenStudy (anonymous):

Find all polar coordinates of point P where P = (1, -pi/6)

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

On the unit circle, \(-\dfrac{\pi}{6}\) is the same angle as \(\dfrac{11\pi}{6}\). Is this what you're looking for?

OpenStudy (anonymous):

well i'm not sure, its saying one of these are the answers A(1, negative pi divided by 6 + (2n + 1)π) or (-1, negative pi divided by 6 + 2nπ) B(1, negative pi divided by 6 + 2nπ) or (-1, negative pi divided by 6 + 2nπ) C(1, negative pi divided by 6 + 2nπ) or (1, pi divided by 6 + (2n + 1)π) D(1, negative pi divided by 6 + 2nπ) or (-1, negative pi divided by 6 + (2n + 1)π)

OpenStudy (anonymous):

You can use the fact I provided to make a general conclusion. On the unit circle, \[\cdots=-\frac{\pi}{6}=\frac{11\pi}{6}=\frac{23\pi}{6}=\cdots=\frac{\pi}{6}+2n\pi=\cdots\] (The angles differ by \(2\pi\), or one full circle about the origin.) That works for all points with \(r=1\): 

OpenStudy (anonymous):

Say you have the point \(\left(-1,-\dfrac{\pi}{6}\right)\):